# MIGE: A Unified Framework for Multimodal Instruction-Based Image Generation and Editing

### v1

@@ 3.3 Joint Training
For subject-driven image generation, we follow the data construction methods of KOSMOS-G Pan et al. (2023) and UNIMO-G Li et al. (2024b), using an LLM to extract entities from captions, which are then fed into Grounded SAM Ren et al. (2024) for segmentation. Subjects200k dataset from OmniControl Tan et al. (2024) is also incorporated for better object preservation. For instruction-based editing, we filter existing datasets and use a rule-based strategy to construct multimodal instructions.
@@
Instruction-based subject-driven image generation is an emerging task that involves two subtasks: instruction-based subject addition and subject replacement. This allows users to add or replace a specified subject in an image using multimodal instructions, as shown in Figure 1. However, there is no sufficient dataset available for this task.
@@
For instruction-based subject addition, we propose a pipeline inspired by SAM-FB He et al. (2024), as shown in Figure 4(a). Starting with the SA-1B Kirillov et al. (2023) dataset, we construct the source image and a multimodal instruction for input-output pairs. We use SAM Kirillov et al. (2023) to segment annotated entities, filter and retain the main subject using MLLM, inpaint the remaining parts to create the background, and then combine the subject name with the target image to generate the multimodal instruction via GPT-4o Hurst et al. (2024). Due to resource limitations, we only process a part of the SA-1B dataset and obtain about 200k samples, but the pipeline can be scaled to generate more. For instruction-based subject replacement, we filter from existing editing data, using Grounded SAM to obtain subject segmentation and construct the multimodal instruction to form input-output pairs, as shown in Figure 4(b). We also introduce virtual try-on data constructed using IDM-VTON Choi et al. (2024), resulting in approximately 110k data samples. More details about training data construction are in Appendix B.
Figure 5: Qualitative comparison for subject-driven image generation (top rows) and instruction-based image editing (bottom rows). We compare the universal model and task-specific models on the two tasks, respectively. The prompts listed in the figure are used for MIGE and vary according to the usage of each model.
4 Experiments
4.1 Implementation Details
@@
MIGE includes a conditional diffusion model and a multimodal encoder. Our design allows flexible selection of various diffusion models, and we initialize with PIXART-$\alpha$ Chen et al. (2023) pretrained at a 512×512 resolution. Parameters introduced for handling conditional inputs are initialized to zero, while the original weights of the model remain unchanged. The multimodal encoder consists of a pretrained Flan-T5-XXL Chung et al. (2024) as LLM for initialization and an image encoding component. This includes query tokens, Q-Former, and a projector, all initialized with a BLIP-2 Li et al. (2023a) checkpoint (pretrain_flant5xxl). The frozen VAE encoder, used as the visual feature extractor, is the same as the one in the diffusion model. Additionally, a zero-initialized MLP layer is introduced in the feature fusing mechanism for progressive visual feature integration.
MIGE is trained on our multi-task dataset using the AdamW optimizer Loshchilov (2017) with a weight decay of 0.03 and a learning rate of 1e-5 for 18 epochs on 48 H20 GPUs, totaling six days of training with a batch size of 960 (20 per GPU). A 1:1 sampling strategy is applied for subject addition and replacement tasks, and during training, there is a 5% chance of dropping either the conditional input or multimodal condition, with an additional 5% chance of dropping both, enabling classifier-free guidance during inference.
4.2 Evaluation Results
@@
As a unified model, MIGE excels in various image generation and editing tasks, outperforming existing task-specific models. In this section, we highlight its strong performance in subject-driven image generation and instruction-based editing, showcasing its emerging capability in instruction-based subject-driven image generation on our new benchmark. For detailed evaluation results, refer to Appendix C. Additional results on these tasks can be found in Figure 12.
4.2.1 Subject-driven Image Generation
Generating images that satisfy both image and text constraints from a multimodal prompt is a challenging task. We compare MIGE with two task-specific methods that also use MLLM to encode multimodal conditions, as well as two universal models, as shown in Table 1. DINO and CLIP-I were used to assess subject fidelity, while CLIP-T evaluated adherence to multimodal instructions. Our results on DreamBench demonstrate that MIGE achieves competitive text fidelity while better preserving subject features, particularly outperforming the DINO metric. The qualitative comparison in Figure 5 further demonstrates that MIGE not only performs better in single-subject generation tasks but also preserves the distinct characteristics of each subject in multi-subject generation, whereas other models either fail to retain all subjects or lose their individual features. This advantage arises from the ability of MIGE to flexibly combine multiple reference entities in multimodal instructions and integrate additional visual features through its feature fusion mechanism.
Methods DINO $\uparrow$ CLIP-I $\uparrow$ CLIP-T $\uparrow$
Task-specific Models
KOSMOS-G Pan et al. (2023) 0.694 0.847 0.287
UNIMO-G Li et al. (2024b) 0.668 0.841 0.329
Universal Models
Omnigen (our test) Xiao et al. (2024b) 0.711 0.800 0.312
Unireal Chen et al. (2024) 0.702 0.806 0.326
MIGE (ours) 0.744 0.830 0.293
          /only_subject data 0.726 0.823 0.289
          /wo_VAE feature 0.741 0.828 0.287
Table 1: Quantitative results for subject-driven image generation on DreamBench. MIGE outperforms universal models in subject preservation and remains competitive with task-specific models.
4.2.2 Instruction-based Image Editing
Instruction-based image editing enables users to modify a source image based on free-form multimodal instructions, including adding, removing, modifying object properties, or altering the overall style. Table 2 presents a quantitative analysis of the Emu Edit and MagicBrush test sets. DINO and CLIP-I evaluate similarity to the source image, while CLIP-T measures alignment with the target description. CLIPdir quantifies the alignment between CLIP text and image embedding changes, whereas L1 and L2 capture pixel-level differences.
As shown in Table 2, MIGE achieves the highest CLIP-T score and surpasses all task-specific models in CLIPdir, demonstrating its superior ability to effectively follow multimodal instructions. As a universal model, MIGE outperforms all other universal models across all metrics on the MagicBrush test set, achieving the lowest L1 and L2 scores and the highest CLIP-I, DINO, and CLIP-T scores, highlighting its strong instruction fidelity and fine-detail preservation. This capability is further illustrated in Figure 5, where MIGE is the only model that accurately follows the instruction to add a Daffy Duck image to the red suitcase without altering other unrelated areas.
@@ 4.2.2 Instruction-based Image Editing
Task-specific Models
InstructPix2Pix Brooks et al. (2023) 0.759 0.831 0.288 0.086 0.122 0.036 0.763 0.843 0.289 0.105 0.097 0.028
MagicBrush Zhang et al. (2024) 0.800 0.857 0.295 0.102 0.085 0.027 0.847 0.888 0.304 0.127 0.062 0.018
UltraEdit Zhao et al. (2024) 0.862 0.875 0.301 0.100 0.050 0.008 0.879 0.897 0.305 0.119 0.042 0.006
Universal Models
PixWizard Lin et al. (2024) 0.824 0.862 0.288 0.066 0.105 0.038 0.865 0.897 0.289 0.054 0.073 0.023
ACE Han et al. (2024) 0.847 0.877 0.299 0.104 0.079 0.028 0.856 0.899 0.304 0.127 0.065 0.023
OmniGen Xiao et al. (2024b) 0.766 0.819 0.299 0.123 0.160 0.065 0.821 0.863 0.289 0.087 0.116 0.045
@@
Table 2: Quantitative results for instruction-based image editing on Emu Edit test set and MagicBrush test set. MIGE achieves the best overall performance on MagicBrush test set among universal models, with strong instruction fidelity and detail preservation.
Figure 6: Qualitative results on the benchmark for the subject addition and subject replacement. The upper section compares subject addition results, while the lower section compares subject replacement. During testing, the <imagehere> placeholder in the multimodal instruction is replaced according to the image sequence. MIGE demonstrates flexibility in editing and excels in subject preservation ability and input-output consistency.
4.2.3 Instruction-based Subject-driven
   Image Editing
Benchmark Construction
Instruction-based subject-driven image editing is a novel task. Existing methods rely on masks or positional coordinates for editing Li et al. (2023b); Yang et al. (2024) but lack support for instruction-based editing. Current benchmarks for subject addition and replacement separately evaluate foreground and background similarities without providing a complete edited image as ground truth, making them unsuitable for this task.
To address these issues, we construct a benchmark of 1,000 test samples: 500 for instruction-based subject addition and 500 for subject replacement. Data is sourced from SEED-Data-Edit Ge et al. (2024), with subjects extracted using Grounded SAM Ren et al. (2024). Captions for target images are generated by GPT-4o and refined through manual review. For compatibility with prior methods, our benchmark also includes masks. Details of the construction process are provided in Appendix A.
Evaluation Results
Our evaluation focuses on editing ability and subject preservation. Editing ability is assessed using DINO, CLIP-I, and CLIP-T: DINO and CLIP-I measure similarity to the ground truth image, while CLIP-T evaluates alignment with the target caption. Subject preservation is evaluated by extracting the edited subject via Grounded SAM and comparing it to the input subject image using DINO and CLIP-I. This separates the evaluation of image-level editing from subject-level feature preservation. Methods that do not support instruction-based editing are marked with a cross in the table, and masks used during testing.
Quantitative comparisons with other methods are shown in Tables 3(a) and 3(b). We computed DINO and CLIP-I metrics between the source and target images, as well as the CLIP-T metric between the source image and target caption, shown in the first row of the table (denoted as “source-target”) as a baseline. In terms of editing ability, MIGE achieves the highest overall improvement across all metrics, demonstrating its effectiveness in performing edits guided by multimodal instructions. As shown in the qualitative comparison in Figure 6, MIGE understands the meaning of “replace” in instructions, rather than simply pasting the entity onto the image. For subject preservation, the results show that MIGE achieves the best performance in both tasks, as demonstrated in Figure 6.
@@ 4.2.3 Instruction-based Subject-driven
Instruction DINO $\uparrow$ CLIP-I $\uparrow$ CLIP-T $\uparrow$ DINO $\uparrow$ CLIP-I $\uparrow$
source-target 0.668 0.842 0.271
PBE Yang et al. (2023) 0.810 0.885 0.304 0.521 0.792
TIGIC Li et al. (2025) 0.789 0.874 0.313 0.453 0.744
MADD He et al. (2024) 0.736 0.852 0.284 0.446 0.742
OmniGen Xiao et al. (2024b) 0.802 0.868 0.296 0.580 0.792
@@
((a)) Results on instruction-based subject replacement.
@@
Instruction DINO $\uparrow$ CLIP-I $\uparrow$ CLIP-T $\uparrow$ DINO $\uparrow$ CLIP-I $\uparrow$
source-target 0.783 0.903 0.295
PBE Yang et al. (2023) 0.843 0.908 0.321 0.495 0.794
TIGIC Li et al. (2025) 0.840 0.901 0.325 0.455 0.753
MADD He et al. (2024) 0.885 0.930 0.316 0.519 0.785
OmniGen Xiao et al. (2024b) 0.791 0.870 0.312 0.605 0.814
@@
((b)) Results on instruction-based subject addition.
Table 3: Quantitative results on instruction-based subject-driven editing. Methods marked with a cross in the instruction column use masks, while the others generate images using multimodal instructions. Overall, MIGE significantly outperforms others in both tasks, showcasing superior editing and subject preservation abilities.
4.3 Further Analysis
4.3.1 Effectiveness of Joint Training
@@
To assess the effectiveness of joint training, we train models separately on individual datasets (denoted as “only_subject data,” “only_edit data,” and “only_compositional data”) and compare their performance with the jointly trained model. The results in Table 1 and Table 2 show that joint training leads to consistent improvements across all metrics, demonstrating that subject-driven generation and instruction-based editing reinforce each other. As shown in Table 3, joint training also enhances the performance of the compositional new task, further highlighting its overall benefits. These findings emphasize both the effectiveness and necessity of joint training. In conclusion, joint training of subject-driven generation and instruction-based editing within our unified framework not only boosts compositional capability but also improves the performance of each individual task.
4.3.2 Effectiveness of Feature Fusing

### v3

@@ 3.3.1. Data Construction
For subject-driven image generation, we follow the data construction methods of KOSMOS-G (Pan et al., 2023) and UNIMO-G (Li et al., 2024c), using an LLM to extract entities from captions, which are then fed into Grounded SAM (Ren et al., 2024) for segmentation. Subjects200k dataset from OmniControl (Tan et al., 2024) is also incorporated for better object preservation. For instruction-based editing, we filter existing datasets and use a rule-based strategy to construct multimodal instructions.
@@
Instruction-based subject-driven image generation is an emerging task involving two subtasks: instruction-based subject addition and subject replacement, which allows users to add or replace a specified subject in an image using multimodal instructions. However, there is no sufficient dataset available for this task.
@@
For instruction-based subject addition, we propose a pipeline inspired by SAM-FB (He et al., 2024a), as shown in Figure 5(a). Starting with the SA-1B (Kirillov et al., 2023) dataset, we construct the source image and a multimodal instruction for input-output pairs. We use SAM (Kirillov et al., 2023) to segment annotated entities, filter and retain the main subject using an MLLM, inpaint the remaining parts to create the background, and then combine the subject name with the target image to generate the multimodal instruction via GPT-4o (Hurst et al., 2024). Due to resource limitations, we process a part of the SA-1B dataset and obtain about 200k samples, but the pipeline can be scaled to generate more. For instruction-based subject replacement, we filter from existing editing data, use Grounded SAM for subject segmentation, and construct multimodal instructions to form input-output pairs, as shown in Figure 5(b). We also introduce virtual try-on data constructed using IDM-VTON (Choi et al., 2024), resulting in approximately 110k samples. More details on training data construction are in Supplementary Material.
4. Experiments
In this section, we present experimental results on three tasks, followed by analysis of the joint training strategy and framework design. For subject-driven image generation, we compare our method with both task-specific models using MLLMs as encoders and recent unified frameworks. For instruction-based image editing, we similarly compare against both task-specific and unified models, including SmartEdit (Huang et al., 2024b) and ml-MGIE (Fu et al., 2023), which also use MLLMs. For instruction-based subject-driven image editing, given the lack of instruction-only baselines, we additionally evaluate mask-dependent methods using masks from MIGEBench.
4.1. Implementation Details
@@
MIGE includes a conditional diffusion model and a multimodal encoder. Our design allows flexible selection of various diffusion models, and we initialize with PIXART-$\alpha$ (Chen et al., 2023) pretrained at a 512×512 resolution. The multimodal encoder consists of a pretrained Flan-T5-XXL (Chung et al., 2024) as LLM for initialization and an image encoding component. This includes query tokens, Q-Former, and a projector, all initialized with a BLIP-2 (Li et al., 2023b) checkpoint (pretrain_flant5xxl). The frozen VAE encoder, used as the visual feature extractor, is the same as the one in the diffusion model. During training, we sample instruction-based subject addition and subject replacement data in a 1:1 ratio. Separately, subject-driven editing and instruction-based editing data are also sampled in a 1:1 ratio. A training strategy is employed with a 5% probability of dropping either the conditional input or the multimodal condition, and an additional 5% chance of dropping both, enabling classifier-free guidance during inference.
4.2. Evaluation Results
@@
As a unified model, MIGE demonstrates exceptional performance across various image generation and editing tasks, even outperforming existing task-specific models. MIGE achieves strong performance in subject-driven generation and instruction-based editing, and shows promising results on instruction-based subject-driven image generation. Detailed results are in Supplementary Material.
Table 1. Quantitative results on the Emu Edit and MagicBrush test sets. MIGE outperforms other universal models on MagicBrush, demonstrating superior instruction alignment and detail preservation.
@@
Task-specific Models InstructPix2Pix (Brooks et al., 2023) 0.759 0.831 0.288 0.086 0.122 0.036 0.763 0.843 0.289 0.105 0.097 0.028
MagicBrush (Zhang et al., 2024) 0.800 0.857 0.295 0.102 0.085 0.027 0.847 0.888 0.304 0.127 0.062 0.018
UltraEdit (Zhao et al., 2024) 0.862 0.875 0.301 0.100 0.050 0.008 0.879 0.897 0.305 0.119 0.042 0.006
SmartEdit (Huang et al., 2024b) 0.752 0.830 0.290 0.100 0.129 0.055 0.821 0.879 0.301 0.122 0.084 0.033
ml-MGIE (Fu et al., 2023) 0.779 0.843 0.276 0.040 0.101 0.034 0.751 0.844 0.274 0.039 0.102 0.034
Unified Models PixWizard (Lin et al., 2024) 0.824 0.862 0.288 0.066 0.105 0.038 0.865 0.897 0.289 0.054 0.073 0.023
ACE (Han et al., 2024) 0.847 0.877 0.299 0.104 0.079 0.028 0.856 0.899 0.304 0.127 0.065 0.023
OmniGen (Xiao et al., 2024) 0.766 0.819 0.299 0.123 0.160 0.065 0.821 0.863 0.289 0.087 0.116 0.045
OneDiff (Le et al., 2024) 0.728 0.819 0.285 0.082 0.136 0.060 0.742 0.840 0.295 0.107 0.128 0.052
@@
4.2.1. Instruction-based Image Editing
Instruction-based image editing allows users to modify a source image using free-form multimodal instructions, such as adding, removing, or altering objects and styles. Table 1 reports results on Emu Edit and MagicBrush (Zhang et al., 2024). DINO and CLIP-I assess similarity to the source image, CLIP-T evaluates alignment with the target caption, CLIPdir measures consistency between text and image embedding changes, and L1/L2 capture pixel-level differences.
As shown in Table 1, MIGE achieves the highest CLIP-T and outperforms all task-specific models on CLIPdir, indicating stronger instruction following. On MagicBrush, it surpasses all universal models across almost all metrics, with the best DINO, CLIP-I, CLIP-T, and lowest L1/L2, demonstrating superior instruction fidelity and detail preservation. This capability is further illustrated in Figure 6, where MIGE is the only model that accurately follows the instruction to add a Daffy Duck image to the red suitcase without altering other unrelated areas. Compared to ml-MGIE, which also uses a MLLM as an encoder, MIGE better preserves input-output consistency and avoids altering irrelevant backgrounds.
Figure 6. Qualitative comparison for instruction-based image editing. MIGE demonstrates superior editing accuracy and instruction understanding.
4.2.2. Subject-driven Image Generation
Generating images that satisfy both image and text constraints from a multimodal prompt is challenging. We compare MIGE with three task-specific methods using MLLMs as encoders and three universal models on DreamBench (Ruiz et al., 2023). DINO and CLIP-I measure subject fidelity, while CLIP-T assesses instruction adherence. As shown in Table 2, MIGE surpasses all unified models in subject preservation ability, as evidenced by the highest DINO and CLIP-I scores, with the DINO score even outperforming task-specific models. Examples in Figure 7 further highlight MIGE’s ability to effectively preserve the identity of each object in multi-object scenarios and combine them based on instructions, demonstrating that our multimodal encoder effectively captures both semantic and visual features simultaneously.
Table 2. Quantitative results for subject-driven image generation on DreamBench. Results with * denote our reproduction. MIGE outperforms universal models in subject preservation and is competitive with task-specific models.
Methods DINO $\uparrow$ CLIP-I $\uparrow$ CLIP-T $\uparrow$
Task-specific Models BLIP-Diffusion (Li et al., 2024a) 0.670 0.805 0.302
KOSMOS-G (Pan et al., 2023) 0.694 0.847 0.287
UNIMO-G (Li et al., 2024c) 0.668 0.841 0.329
Unified Models Omnigen * (Xiao et al., 2024) 0.711 0.800 0.312
UniReal (Chen et al., 2024c) 0.702 0.806 0.326
OneDiffusion * (Le et al., 2024) 0.582 0.750 0.240
MIGE (ours) 0.744 0.830 0.293
/only_subject data 0.726 0.823 0.289
/wo_VAE feature 0.741 0.828 0.287
Figure 7. Qualitative comparison for subject-driven image generation. MIGE exhibits superior subject preservation, even in multi-subject scenarios.
Table 3. Quantitative results on instruction-based subject-driven editing. Methods marked with a cross in the instruction column use masks, while the others generate images using multimodal instructions. Overall, MIGE significantly outperforms others in both tasks, showcasing superior editing and subject preservation abilities.
(a) Results on instruction-based subject replacement.
@@ 4.2.2. Subject-driven Image Generation
DINO $\uparrow$ CLIP-I $\uparrow$ CLIP-T $\uparrow$ DINO $\uparrow$ CLIP-I $\uparrow$
I-O sim 0.668 0.842 0.271
Mask-denpendent Models PBE (Yang et al., 2023) 0.810 0.885 0.304 0.521 0.792
AnyDoor (Chen et al., 2024b) 0.817 0.890 0.307 0.551 0.789
TIGIC (Li et al., 2025c) 0.789 0.874 0.313 0.453 0.744
MADD (He et al., 2024a) 0.736 0.852 0.284 0.446 0.742
Mask-free Models OmniGen (Xiao et al., 2024) 0.802 0.868 0.296 0.580 0.792
@@
/wo_VAE feature 0.852 0.900 0.308 0.633 0.819
(b) Results on instruction-based subject addition.
@@
DINO $\uparrow$ CLIP-I $\uparrow$ CLIP-T $\uparrow$ DINO $\uparrow$ CLIP-I $\uparrow$
I-O sim 0.783 0.903 0.295
Mask-denpendent Models PBE (Yang et al., 2023) 0.843 0.908 0.321 0.495 0.794
AnyDoor (Chen et al., 2024b) 0.863 0.930 0.324 0.533 0.799
TIGIC (Li et al., 2025c) 0.840 0.901 0.325 0.455 0.753
MADD (He et al., 2024a) 0.885 0.930 0.316 0.519 0.785
Mask-free Models OmniGen (Xiao et al., 2024) 0.791 0.870 0.312 0.605 0.814
@@
/wo_VAE feature 0.902 0.934 0.322 0.633 0.831
4.2.3. Instruction-based Subject-driven Image Editing
Benchmark Construction
Instruction-based subject-driven image editing is a novel task. Existing methods rely on masks or positional cues for editing (Li et al., 2023a; Yang et al., 2024) but lack support for instruction-only editing. The AnyEdit-Test Benchmark (Yu et al., 2024) supports few instruction templates and evaluates only one type of object editing based on automatically extracted masks. Current benchmarks (Yang et al., 2023; Chen et al., 2024b) for subject addition and replacement separately evaluate foreground and background similarities without providing a complete edited image as ground truth, making them unsuitable for this task.
To address these issues, we manual construct MIGEBench, a benchmark consisting of 1,000 test samples: 500 for instruction-based subject addition and 500 for subject replacement. More details can be found in the Supplementary Material.
Evaluation Results
Our evaluation focuses on editing ability and subject preservation. DINO and CLIP-I measure the similarity of the generated image to the ground truth, while CLIP-T evaluates alignment with the target caption, indicating editing ability. Subject preservation is evaluated by extracting the edited subject via Grounded SAM (Ren et al., 2024) and comparing it to the input subject image using DINO and CLIP-I. This separates the evaluation of image-level editing from subject-level feature preservation. Methods that do not support instruction-based editing use masks provided in MIGEBench during testing.
Quantitative comparisons with other methods are shown in Tables 3(a) and 3(b). We compute the DINO, CLIP-I, and CLIP-T scores between the input image and the ground truth. The first row of the table (denoted as ”I-O sim”) provides a baseline for evaluating the effectiveness of editing ability. Mask-dependent models can easily preserve the edit-unrelated regions, but MIGE still outperforms all methods, highlighting the effectiveness of conditional inputs and MIGE’s precise understanding of instructions. The CLIP-T score measures instruction adherence in mask-free models. MIGE significantly outperforms OmniGen in both tasks, demonstrating the advantage of our unified approach. MIGE also achieves the best performance on subject preservation in both tasks. As shown in Figure 8, MIGE successfully understands the task of replacing a bird with an owl, rather than simply overlaying or adding an owl, demonstrating the effectiveness of the multimodal encoder which simultaneously encoding the semantic and visual information of the reference image.
Figure 8. Qualitative results on the benchmark for the subject addition and subject replacement. The upper section compares subject addition results, while the lower section compares subject replacement. During testing, the ¡imagehere¿ placeholder in the multimodal instruction is replaced according to the image sequence. MIGE demonstrates flexibility in editing and excels in subject preservation ability and input-output consistency.
4.3. Further Analysis
4.3.1. Effectiveness of Joint Training
@@
To assess the effectiveness of joint training, we train models separately on individual datasets (denoted as “only_subject data,” “only_edit data,” and “only_compositional data”) and compare their performance with the jointly trained model. The results in Tables 1, 2 and 3 show that joint training leads to consistent improvements across all metrics, demonstrating that subject-driven generation and instruction-based editing reinforce each other. As shown in Table 3, joint training also enhances the performance of the compositional new task, further highlighting its overall benefits. These findings emphasize both the effectiveness and necessity of joint training. In conclusion, joint training of subject-driven generation and instruction-based editing within our unified framework not only boosts compositional capability but also improves the performance of each individual task. This validates our motivation and provides further theoretical support for unified models for multiple tasks.
4.3.2. Effectiveness of Feature Fusing